OpenStudy (anonymous):

Solve the logarithmic equation. Round to the nearest ten-thousandth if necessary. \[ 3 \log (2x) = 4 \]

OpenStudy (anonymous):

3 log(2x) = 4 log(2x) = 4/3 2x = 10^(4/3) x = (1/2) 10^(4/3)

OpenStudy (anonymous):

What log properties did you use?

OpenStudy (jdoe0001):

\(\bf 3log(2x)=4\implies 3log_{10}(2x)=4\implies log_{10}(2x)=\cfrac{4}{3}\\ \quad \\ \textit{log cancellation rule of }\qquad \Large \color{red}{a^{log_ax}=x}\qquad thus\\ \quad \\ log_{10}(2x)=\cfrac{4}{3}\implies \Large 10^{log_{10}(2x)}=10^{\frac{4}{3}}\implies 2x=10^{\frac{4}{3}}\)